OpenStudy (anonymous):

Solve the equation by completing the square. round to the nearest hundredth if necessary. x^2 + 2x = 3

OpenStudy (anonymous):

what is half of 2?

OpenStudy (anonymous):

1.5 @satellite73

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[ (x+d)^2=x^2+2dx+d^2 \]

OpenStudy (anonymous):

half of two?

OpenStudy (anonymous):

\[ 2d=2 \]

OpenStudy (anonymous):

the half of 2 is 1 sorry

OpenStudy (anonymous):

you dont have to complete the square (x-1)(x+3)

OpenStudy (anonymous):

ok then start with \[(x+1)2=3+?\] now we fill in the ? what is one squared?

OpenStudy (anonymous):

one squared is 1

OpenStudy (anonymous):

ok so we go to \[(x+1)^2=3+1\] or \[(x+1)^2=4\]

OpenStudy (anonymous):

then taking the square root, we get \[x+1=-2\] or \[x+1=2\] solve the first one, get \[x=-3\]and the second gives \[x=1\]

OpenStudy (anonymous):

ok i got that

OpenStudy (anonymous):

but is that the complete answer?

OpenStudy (anonymous):

@satellite73